OpenStudy (anonymous):

can someone show the steps and answer to this problem i did it wrong and im confused (x + 6)(x − 6) and (x − 6)(x − 6)

OpenStudy (diyadiya):

(x+6)(x-6) you can use the identity \[(a+b)(a-b)=a^2-b^2 \]

OpenStudy (diyadiya):

here a =x & b= 6

OpenStudy (diyadiya):

Just substitute the values :)

OpenStudy (diyadiya):

can you try that?

OpenStudy (anonymous):

(x+6)(x-6)=x^2-6x+6x-36=x^2-36 (x-6)(x-6)=x^2-12x+36